Salza Follow the gudeline and fill all the missing numbers for cash flow identity analysis. Based on your ca culation, complete the following multiple choice questions. Cash flow identity Salza Technology Corporation increased its sales from $375,000 in 2012 to $450,000 in year 2013 as is shown in the firms income statements presented below. LeAnn Sands, chief executive ohicer (CEO) and founder of the firm expressed concern that the cash account and ihe finms marketable securities declined substantially between 2012 and

Quiz: Homework2 2013. Salzas complete balance sheets are also shown below, Ms. Sands is soeking your assistance in the preparation of a statement of cash flows for Salza Technology Hint: use my slides for reference Page Income statement (in S Thousands) 2013 2012 450.00 375.00 270.00 225.00 180.00 150.00 46.00 46.00 30.00 25.00 104.00 79.00 4.00 4.00 100.00 75.00 30.00 20.00 70.00 55.00 20.00 17.00 50.00 38.00 Net sales Less: cogs Gross profit Less operating exp Less. Depr EBIT Less Interest Income before tax Less: Income taxes Net income Cash dividend Addition to retained earnings Balance sheet (in $ Thousands) Cash Account receivable Inventories Current asset Gross fixed asset less accumuled depr Net fixed asset Total assets Account payable Bank loarn Accrued liabilities Total current liabilities Long-term debt Total liabilities Common stock Retained earnings Total equity 2013 2012 16.00 39.00 80.0050.00 204.00 151.00 300.00 24000 290.00 200.00 125.00 95.00 165.00 105.00 465.00 345.00 45.00 30.00 27.00 20.00 23.00 10.00 9500 80.00 15.00 15.00 10.00 75.00 120 0085.00 23500 185 00 355.00 270 0 ps:/lynn.instructure.com/courses/6822/quizzes 40302/take

Salza: How much is Operating cash flow? O 100.00 O 200.00 0 104.00 0 204.00 Question 32 Salza: How much is Net Capital Spending (NCS)? ○ 120 060 090 030

300 Question 34 0 32 o 200 0 35 Question 35
Answer #1
A B C D E F G H I J K L
2
3 31)
4
5 Operating Cash Flow in 2013 =EBIT - Taxes +depreciation
6 =104 - 30 + 30
7 $104.00
8
9 Hence operating cash flow is $104.00
10 Thus the third option is correct.
11
12 32)
13 Net Capital Spending in 2013 =Gross Fixed Asset in 2013 - Gross Fixed asset in 2012
14 =290-200
15 90
16
17 Hence Net Capital Spending is 90
18 Thus the third option is correct.
19
20 33)
21
22 Net Operating working capital in 2013 = Oprating Current Assets - Operating Current Liabilities
23 =Cash + Account Receivable + Inventories - (Accounts Payable + Accrued Liabilities)
24 =16+80+204-(45+23)
25 =300-68
26 232
27
28 Hence Net Operating working capital in 2013 232
29 Thus the second option is correct.
30
31 34)
32 Net Operating working capital in 2012 = Oprating Current Assets - Operating Current Liabilities
33 =Cash + Account Receivable + Inventories - (Accounts Payable + Accrued Liabilities)
34 =39+50+151-(30+10)
35 =240-40
36 200
37 Net Operating working capital in 2013 232
38
39
40 Change in net Operating working capital in 2013 =Net operating working capital in 2013 - Net Operating capital in 2012
41 =232-200
42 32
43
44 Change in net Operating working capital in 2013 32
45 Hence the first option is correct.
46
47 35)
48
49 Cash Flow from Asset =Operating Cash Flow - Change in Net operating working capital - Net Capital Spending
50 =104-32-90
51 ($18.00) =D9-D44-D17
52
53 Hence cash flow from asset is ($18.00)
54 Thus the third option is correct.
55